OpenStudy (anonymous):

Simplify the given equation. 6 - (3x + 10) + 4(2 - x) = 15

OpenStudy (anonymous):

remove bracket 6-(3x+10) + 4(2-x)=15 6-3x-10 + 4*2 -4x =15 -7x +4=15 -7x =11 x= -11/7

OpenStudy (anonymous):

u should try at least